OpenStudy (anonymous):

Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of (f − g)(−2)? −8 −5 −11 3

OpenStudy (anonymous):

I'm guessing you mean when x = -2? f(-2) = 3x − 2 = -8 g(-2) = x2 − 1 = 3 f - g = -8 -3 = -11 -11 is your answer!